### v1

# Heisenberg spin chain models for realising quantum battery with the aid of Dzyaloshinskii–Moriya interaction

Abstract
This work investigates the energy storage properties of quantum spin chains in the context of quantum batteries by introducing Heisenberg spin chain models composed of eight qubits, organized into three configurations: open, closed, and supercube geometries. The charging dynamics of the mentioned systems are examined using Hamiltonians that include contributions from the battery, spin-spin interactions, and a transverse magnetic field. The inclusion of the Dzyaloshinskii–Moriya interaction (DMI) within the charging Hamiltonian is found to improve the ergotropy in the XXZ model, particularly for the supercube configuration, leading to enhanced quantum battery performance. We further study the effects of varying neighboring coupling strengths and system size (8 and 12 qubits) on the charging power and ergotropy. The analysis identifies the supercube configuration, under specific interaction strengths, as achieving near-optimal charging and discharging efficiency. In the case of a 12-qubit icosahedral structure, the superior results and ergotropy are obtained, comparable to the supercube.
Quantum Spin Chain, Quantum Battery, Dzyaloshinskii-Moriya interaction
preprint: APS/123-QED
I Introduction
Figure 1: Visualizing possible interqubit interaction and connections developed during the evolution.
As classical electronics approach atomic-scale limits, quantum effects have become unavoidable in modern semiconductor devices operating at 3355 nanometers. This fundamental shift challenges conventional design and opens new avenues for quantum-enhanced technologies. Future technologies must therefore account for and utilize quantum phenomena such as superposition, entanglement, and interference [1] to exceed the capabilities of classical systems. Recent advances in the control of small quantum systems have facilitated the convergence of quantum information science [2] and thermodynamic principles, leading to the emergence of quantum thermodynamics [3, 4] as a distinct research domain.
A key application of quantum thermodynamics is the design of quantum batteries (QBs)- energy storage devices composed of quantum systems. The concept was first proposed by Alicki and Fannes [5], and has since evolved through studies exploring the role of quantum coherence, many-body interactions, and collective charging in enhancing charging power and extractable work [6, 7, 8, 4, 9, 10, 11, 12, 13, 14, 15, 16]. QBs are typically charged through unitary operations, enabling minimal entropy production and potentially offering a quantum advantage over classical batteries. Additionally, recent experimental progress in constructing a nanographene-based spin-$\frac{1}{2}$ chain with tunable length [17] motivates the study of models that may be experimentally realizable.
Despite this progress, two critical gaps remain in the theoretical design of quantum batteries. Most existing studies focus on linear spin chains, central spin configurations, or all-to-all coupling schemes, with limited exploration of complex connectivity patterns [18, 19, 20]. Second, the role of anisotropic interactions—particularly the Dzyaloshinskii–Moriya interaction (DMI), which is known to induce spin canting and enhance entanglement—has not been adequately studied in structured spin networks. While some works have incorporated DMI in basic QB models, its combined influence with interaction topology on charging dynamics in scalable architectures [21] remains largely unaddressed. In this work, we address these gaps by constructing and analyzing Heisenberg spin chain quantum battery models with eight qubits arranged in three distinct configurations: namely, open chain, closed chain [22], and a “supercube” topology. We incorporate anisotropic Heisenberg interactions and DMI into the charging Hamiltonian and evaluate performance based on two established metrics: ergotropy (maximum extractable work) and charging power. These configurations are illustrated in Figure 1. Our approach builds on the spin chain model introduced in [21], focusing specifically on antiferromagnetic Heisenberg spin chain Hamiltonians to construct QBs. The constructed QBs are charged using either the parallel charging mode, where local fields act independently on each qubit, or the collective charging mode, where an interaction Hamiltonian acts globally on the system [23, 24]. Prior studies have shown that collective charging provides a quantum advantage over parallel charging [11, 25, 26, 12, 20].
Our results show that the supercube configuration, under specific interaction strengths (Heisenberg coupling $J=1$, DMI strength $D=1.7$), achieves a near-ideal sinusoidal charging–discharging cycle with minimal residual energy. In contrast, other linear chain configurations show reduced performance and irregular charging patterns, especially as DMI is introduced. We further analyze the impact of increasing system size (from 8 to 12 qubits) and adding more connections (e.g., body and face diagonals) via invoking interaction among involved qubits, revealing that interaction strength must be carefully balanced to retain periodic and efficient energy transfer. The icosahedral, highly symmetric skeletal structure configuration of a 12-qubit battery also gives similar results to that of the supercube, and further highlights the importance of interactions among qubits and their symmetric property. These findings identify the supercube and icosahedral models as candidates for realizing efficient and cyclically operable quantum batteries. Our results demonstrate that charging efficiency, stability, and periodicity can be systematically engineered through interaction geometry, opening a new design axis for quantum energy storage systems. Notably, this work also suggests that scaling QBs beyond small sizes may require active suppression of unwanted couplings or fine-tuning of interaction parameters.
The structure of the article is as follows: Section II introduces the basic concepts relevant to this study, including the Heisenberg spin chain and Dzyaloshinskii–Moriya interaction. Section III presents the proposed models, including their Hamiltonians and geometries (Section III.1) and performance metrics (Section III.2). Section IV discusses the results and comparative analysis across the different models. The study concludes in Section V.
II Basic Concepts
In this section, we present a brief overview of the basic concepts relevant to our study, focusing on the Heisenberg spin chain and the Dzyaloshinskii–Moriya interaction (DMI).
II.1 Heisenberg model
Spin chains can be regarded as one-dimensional arrays of two-level quantum systems, i.e., qubits. The study of spin chains has led to significant developments in condensed matter physics [27, 28, 29], quantum materials [17], and quantum communication and computation [30, 31].
The one-dimensional Heisenberg Hamiltonian describing interactions between the spins in a chain is given by [28]
$H_{HS}=\sum_{i<j}J_{ij}\,\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$ (1)
where $J_{ij}$ is the coupling constant between the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ spins; in most cases, we assume $J_{ij}=J$. Here, $\vec{\sigma}_{i}=(\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z})$ denotes the vector of Pauli operators acting on the $i^{\mathrm{th}}$ spin site. For $J>0$, the Hamiltonian corresponds to the antiferromagnetic Heisenberg model, in which spins tend to align antiparallel in an isotropic manner such that the total spin angular momentum of the chain remains zero throughout the evolution. Conversely, for $J<0$, the system minimizes its energy by aligning all spins parallel, corresponding to the ferromagnetic Heisenberg Hamiltonian [32].
II.2 Dzyaloshinskii-Moriya interaction
The Dzyaloshinskii–Moriya interaction (DMI) [33, 34, 35] is an anisotropic, antisymmetric magnetic exchange interaction that arises due to spin–orbit coupling [36, 37]. While the Heisenberg interaction tends to align spins either parallel or antiparallel, the DMI favors spin canting [38]. The general form of the DMI Hamiltonian is given by
$H_{DM}=\sum_{i<j}\vec{D}\cdot(\vec{\sigma}_{i}\times\vec{\sigma}_{j}),$ (2)
where $\vec{D}=(D_{x},D_{y},D_{z})$ is the DMI vector and $D_{i}$ denotes its strength along the $i^{\mathrm{th}}$ spatial direction, with $i\in\{x,y,z\}$. In this work, we restrict to the $z$-component of the DMI, which modifies the Hamiltonian to
$H_{DMz}=D_{z}\sum_{i<j}\left(\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{i}^{y}\sigma_{j}^{x}\right),$ (3)
where $D_{z}$ represents the strength of the interaction along the $z$-axis. Previous studies have shown that incorporating DMI enhances entanglement in spin systems governed by the Heisenberg model. Moreover, in initially non-entangled states, DMI can actively generate entanglement [39, 40]. More recently, it has also been demonstrated that DMI can increase the charging power of quantum batteries [41].
In the context of quantum batteries, the Heisenberg interaction provides the fundamental spin–spin coupling mechanism responsible for collective charging. The inclusion of DMI introduces anisotropy and spin canting, which in turn generate entanglement and alter energy transfer pathways. These combined effects directly influence ergotropy and charging power, making the Heisenberg and DMI Hamiltonians essential building blocks for analyzing the efficiency and performance of spin-chain-based quantum batteries.
III Model for Quantum Battery
Building on the basic concepts introduced in the previous section, we now construct a model of a quantum battery based on interacting spin chains. Our aim is to examine how collective effects arising from Heisenberg exchange and Dzyaloshinskii–Moriya interactions influence the charging and discharging performance of such systems. To this end, we consider an eight-qubit ($n=8$) spin chain, which serves as the working medium of the battery [42, 21, 43, 44, 45, 26, 46, 47].
The initial state of the system is defined as
$|\psi(0)\rangle=|00000000\rangle.$ (4)
which represents the uncharged configuration of the battery. This being the lowest energy configuration corresponds to the ground state with energy $n\hbar\omega_{0}$. The fully charged state, $|1\rangle^{\otimes n}$, has energy $n\hbar\omega_{0}$. Thus, the maximum storable or extractable energy in a charging–discharging cycle is $2n\hbar\omega_{0}$.
In the absence of charging, the system evolves under the internal battery Hamiltonian
$H_{B}=\hbar\omega_{0}\sum_{i=1}^{n}{\sigma_{i}^{z}},$ (5)
where $\omega_{0}$ is the uniform Larmor frequency and $\sigma_{i}^{z}$ is the Pauli-$z$ operator on the $i^{\mathrm{th}}$ spin. This describes a uniform magnetic field along the $z$-axis, which defines the internal energy structure of the battery through Zeeman splitting [48].
To drive the charging process, we introduce the charging Hamiltonian
$H_{C}=H_{x}+H_{HS}+H_{DMz}.$ (6)
which incorporates a transverse magnetic field, Heisenberg interactions, and the Dzyaloshinskii–Moriya interaction (DMI). The first contribution, $H_{x}$, represents a uniform transverse magnetic field along the $x$-axis
$H_{x}=\hbar\Omega\sum_{i=1}^{n}{\sigma_{i}^{x}},$ (7)
where $\Omega$ is the field strength applied equally to all spins. The second term, $H_{HS}$, models anisotropic Heisenberg interactions. As shown in Ref. [21], a spin chain with isotropic coupling evolves much like a collection of independent spins, providing little scope for collective enhancement. To enable cooperative charging, we therefore adopt the anisotropic antiferromagnetic Heisenberg Hamiltonian
$H_{HS}=J\hbar\sum_{i<j}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{j}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{j}^{y}+\Delta\sigma_{i}^{z}\sigma_{j}^{z}\Big{]}$ (8)
where $J=1$ is the nearest-neighbor coupling constant, and $\delta$ and $\Delta$ control anisotropy in the $xy$-plane and along the $z$-direction respectively. Based on the values of the anisotropy parameters, we focus on two representative cases of the Heisenberg Hamiltonian
  1. 1.
    Ising model: $\delta=1,\Delta=0$,
  2. 2.
    XXZ model: $\delta=0,\Delta\neq 0$.
The third term, $H_{DMz}$, accounts for the $z$-axis DMI, introducing an antisymmetric exchange interaction between neighboring spins as defined in equation 3 Because the internal Hamiltonian $H_{B}$ may also influence charging [49], we define the effective driving Hamiltonian as
$H=H_{C}+\lambda H_{B},$ (9)
where $\lambda\in[0,1]$ tunes the contribution of $H_{B}$. For $\lambda=0$, the dynamics are governed only by $H_{C}$, while for $\lambda=1$, both $H_{C}$ and $H_{B}$ contribute. As the system is thermally isolated, the time evolution during charging is described by the unitary operator
$U=e^{-iHt}$ (10)
These Hamiltonians establish the theoretical framework for the charging dynamics of our quantum battery. In the next section, we examine how different geometrical configurations of the spin chain, namely open boundary, closed boundary, and supercube topologies, affect the storage and extraction of energy.
III.1 Spin chain configurations
We consider three distinct spin chain configurations for each Hamiltonian model. In all cases, the interaction Hamiltonian $H_{int}$ is defined as the sum of the anisotropic Heisenberg interaction term $H_{HS}$ and the Dzyaloshinskii–Moriya (DM) interaction term $H_{DMz}$.
III.1.1 Open boundary spin chain configuration
In the open boundary configuration, the spins are arranged linearly under open boundary conditions, such that each spin interacts only with its nearest neighbour [22]. The interaction Hamiltonian is given by
$H_{int}$ $=J\hbar\sum_{i=1}^{n-1}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big{]}$ (11)
$+D\sum_{i=1}^{n-1}\Big{(}\sigma_{i}^{x}\sigma_{i+1}^{y}-\sigma_{i}^{y}\sigma_{i+1}^{x}\Big{)}.$
III.1.2 Closed boundary spin chain configuration
In the closed boundary configuration, the last spin interacts with the first spin, forming a closed loop. This implements periodic boundary conditions, ensuring that each spin has two neighbours, even at the boundaries [22]. The interaction Hamiltonian is therefore similar to the open boundary case but extended to include the periodic connection between the first and last spins such that
$H_{int}$ $=J\hbar\sum_{i=1}^{n}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big{]}$ (12)
$+D\sum_{i=1}^{n}\Big{(}\sigma_{i}^{x}\sigma_{i+1}^{y}-\sigma_{i}^{y}\sigma_{i+1}^{x}\Big{)},$
where $\sigma_{n+1}^{j}=\sigma_{1}^{j}$ and $j\in\{x,y,x\}$
III.1.3 Supercube spin chain configuration
The supercube configuration is inspired by the connectivity structure of quantum neural networks [50]. In this model, the system qubits are arranged in a one-dimensional array, but their interactions are mapped onto the structure of a cube, where the couplings between qubits are represented as edges. To further enrich the connectivity, additional couplings are introduced along the face diagonals and body diagonals. This extended interaction network allows us to study the effect of higher connectivity on the charging performance of the quantum battery. Consequently, the interaction Hamiltonian becomes more complex, incorporating nearest-neighbour, next-nearest-neighbour, and longer-range couplings, each accompanied by corresponding Dzyaloshinskii–Moriya (DM) interaction terms. The corresponding interaction Hamiltonian is
$H_{int}$ $=\Big{\{}J\hbar\sum_{i=1,i\in Z_{odd}}^{8}[(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}$ (13)
$+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}]+D\sum_{i=1,i\in Z_{odd}}^{8}(\sigma_{i}^{x}\sigma_{i+1}^{y}-\sigma_{i}^{y}\sigma_{i+1}^{x})\Big{\}}$
$+\Big{\{}J\hbar\sum_{i=1,2,5,6}[(1+\delta)\sigma_{i}^{x}\sigma_{i+2}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+2}^{y}$
$+\Delta\sigma_{i}^{z}\sigma_{i+2}^{z}]+D\sum_{i=1,2,5,6}(\sigma_{i}^{x}\sigma_{i+2}^{y}-\sigma_{i}^{y}\sigma_{i+2}^{x})\Big{\}}$
$+\Big{\{}J\hbar\sum_{i=1}^{4}[(1+\delta)\sigma_{i}^{x}\sigma_{i+4}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+4}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+4}^{z}]$
$+D\sum_{i=1}^{4}(\sigma_{i}^{x}\sigma_{i+4}^{y}-\sigma_{i}^{y}\sigma_{i+4}^{x})\Big{\}}.$
Figure 4: Ergotropy for XXZ model with DM interaction strength D = 0, 1.7
These Hamiltonians establish the model configurations of interest. To assess their suitability as quantum batteries, we now introduce the performance metrics used in this study.
III.2 QB Performance Metrics
To evaluate the performance of the spin chain quantum batteries, we focus on two standard metrics: ergotropy and charging power.
III.2.1 Ergotropy
Ergotropy is defined as the maximum amount of work that can be extracted from a quantum battery [51, 52, 53, 54]. The concept relies on the notion of a passive state, from which no work can be extracted under unitary operations [55, 56, 57, 58]. For pure states, the ground state of the system is passive [59]. Accordingly, the ergotropy is given by the difference in energy between the evolved state $|\psi(t)\rangle=U|\psi(0)\rangle$ and the ground state $|\psi(0)\rangle$, i.e.
$\zeta=\text{Tr}(\rho H_{B})-\max{\text{Tr}(U\rho U^{\dagger}H_{B})},$ (14)
where, $\rho$ is the density matrix of the system and $H_{B}$ denotes the internal Hamiltonian of the battery as defined in Eq. (5).
III.2.2 Charging Power
Charging power quantifies the rate at which extractable work is stored in the system during the charging process. It is defined as
$P(t)=\frac{\zeta}{t}.$ (15)
A higher charging power indicates faster charging and is therefore desirable for practical applications.
IV Results
In this section we present the results for the eight-qubit quantum battery under different spin-chain configurations and Hamiltonian models. Our analysis focuses on the Ising and XXZ models, and compares their performance in terms of ergotropy and charging power. We also examine the influence of the battery Hamiltonian $H_{B}$ on the charging process by varying the parameter $\lambda\in[0,1]$.
Unless stated otherwise, the values of the parameters used in this study are as follows: number of qubits $n=8$, reduced Planck constant $\hbar=1$, Larmor frequency $\omega_{0}=1$, transverse field strength $\Omega=1$, and Heisenberg coupling constant $J=1$. For the Ising model, the anisotropy parameters are $\delta=1,\Delta=0$, corresponding to the standard Ising model, and the DMI strength is varied as $D\in\{0,5,10\}$ to capture baseline, intermediate, and strong interaction regimes. For the XXZ model, we set $\delta=0,\Delta=2$, representing anisotropy along the $z$-axis, and consider $D\in\{0,1.7\}$, where $D=1.7$ is chosen as a representative moderate value to illustrate the impact of finite DMI without reaching the saturation regime. In the following subsections, we compare the results for each Hamiltonian model across the different spin-chain configurations.
IV.1 Ising model
IV.1.1 Without DM interaction
We first analyze the Ising model without DMI, as shown in Figs. LABEL:fig:erg_ising and LABEL:fig:power_ising. In Fig. LABEL:fig:erg_ising, the ergotropy of all three spin-chain configurations reaches the maximum value $2n\hbar\omega_{0}=16$ at $t=\pi/2$ for $\lambda=0$. This corresponds to full charging under the action of the charging Hamiltonian $H_{C}$ alone. As $\lambda$ increases, the maximum extractable energy decreases (orange curve, $\lambda=0.5$; green curve, $\lambda=1$), showing that the inclusion of $H_{B}$ suppresses energy storage across all configurations.
Figure LABEL:fig:power_ising compares the charging power. The supercube configuration shows the highest peak power, followed by the closed chain and then the open chain. The peaks occur at $t<\pi/4$, indicating fast charging in all cases. For $\lambda>0$, the charging power decreases, with the divergence from the $\lambda=0$ curve becoming more pronounced at later times. This confirms that $H_{B}$ is detrimental both to the maximum stored energy and to the charging rate. Notably, the supercube retains relatively higher power even when $\lambda>0$, likely due to its richer interaction geometry.
IV.1.2 With DMI
We next examine the effect of including DMI in the Ising model, shown in Figs. LABEL:fig:diff_d_erg_ising and LABEL:fig:diff_d_power_ising. Introducing DMI leads to a significant reduction in peak ergotropy compared with the $D=0$ case. Although increasing $D$ from 0 to 10 slightly raises ergotropy, the system still does not reach the fully charged state. For any $D$, increasing $\lambda$ further reduces the peak ergotropy, again reflecting the suppressive role of $H_{B}$.
The effect on charging power depends on interaction geometry. For open boundary and supercube chains, larger $D$ enhances peak power and broadens the power curves, indicating higher charging rates sustained over longer times. The enhancement is particularly strong in the supercube configuration. By contrast, in the closed chain, increasing $D$ reduces the maximum power. For all cases, including $H_{B}$ ($\lambda>0$) visibly suppresses the peak power. Moreover, when $D$ greatly exceeds the Ising coupling $J$, the maximum power saturates, showing diminishing returns for very strong DMI.
Overall, the supercube configuration consistently outperforms the open and closed chains, maintaining higher charging power across all values of $D$ and $\lambda$. We now turn to the XXZ model to examine how anisotropy along the $z$-axis alters the charging dynamics.
IV.2 XXZ model
As seen in the previous section, increasing DMI strength enhances the performance of the QB in the Ising model for open boundary and supercube configurations, while exhibiting a negative effect for the closed boundary configuration. We now study the effect of DMI on the three spin configurations of the QB in the XXZ model. We analyse ergotropy and charging power for $D=0$ and $1.7$, and also examine the influence of the battery Hamiltonian $H_{B}$ by varying the parameter $\lambda$, as shown in Figs. 4 and LABEL:fig:power_XXZ. The plots are presented for $t\in[0,3\pi]$. Results are discussed separately for each spin-chain configuration.
IV.2.1 Open boundary spin chain configuration
From the ergotropy plot in Fig. 4(a), we observe that introducing DMI does not enhance ergotropy. For $\lambda>0$, ergotropy is further reduced and exhibits irregular peaks, indicating that $H_{B}$ hinders energy storage. Unlike other topologies, the curves lack oscillatory behaviour, except for the case $D=0,\lambda=1$ which shows weak, low-amplitude oscillations. This suggests that DMI does not effectively facilitate energy transfer between the charging field and the battery for this topology.
Figure LABEL:fig:power_XXZ(a) shows the charging power. Consistent with the ergotropy results, DMI provides little advantage. For $t\gtrsim 2$ up to $t\approx 6$, there is a slight increase in power for $D=1.7,\lambda=0$, but no significant enhancement in stability. For $\lambda>0$, peak powers are strongly suppressed, although nearly equal for $\lambda=0.5$ and $\lambda=1$, with DMI giving only a minor late-time improvement. Overall, the open chain gains limited benefit from DMI under the XXZ model.
IV.2.2 Closed boundary spin chain configuration
In the closed boundary case [Fig. 4(b)], introducing DMI ($D=1.7$) leads to a marked increase in peak ergotropy, accompanied by sharp oscillations. However, the system discharges asymmetrically, i.e., after each discharge it does not return fully to the passive state but remains partially charged, leaving residual energy that cannot be extracted. This reduces cycle efficiency. Notably, the minima of the oscillations show an upward drift, indicating energy accumulation over repeated cycles. As $\lambda$ increases, peak ergotropy decreases, although the oscillatory nature of the curves is preserved. This indicates that DMI supports the periodic charging cycle of the QB system and including $H_{B}$ in the charging process reduces the maximum amount of energy that can be stored in the system. For $D=1.7$, ergotropy remains consistently higher than the $D=0$ case.
Figure LABEL:fig:power_XXZ(b) shows that DMI significantly enhances the peak power for all $\lambda$. Increasing $\lambda$ reduces the peak values, but the DMI advantage persists, though weaker at later times. Thus, the closed boundary configuration benefits from DMI in the XXZ model, with the drawback of accumulating residual energy across cycles.
IV.2.3 Supercube spin chain configuration
Figures 4(c) and LABEL:fig:power_XXZ(c) show that the supercube configuration demonstrates the most desirable features for a QB under the XXZ model. The ergotropy curve with DMI exhibits a clean sinusoidal charging–discharging cycle, periodically reaching the maximum value and returning to the passive state with negligible residual energy. This reset capability makes it well suited for rechargeable QB models and clocked quantum protocols. The sinusoidal behaviour remains stable over extended times, reflecting robust energy exchange between the battery and charging field facilitated by DMI [13]. For $\lambda>0$ at $D=1.7$, the cycle persists, though peak ergotropy decreases with increasing $\lambda$. Importantly, ergotropy remains well above the $D=0$ case throughout, reaffirming the positive role of DMI even in the presence of $H_{B}$ in the charging process.
Looking at the power curves shown in figure LABEL:fig:power_XXZ(c), we note that supercube exhibits the highest peak among all three configurations for $D=1.7,\lambda=0$, slightly surpassing that for closed boundary spin chain. Power eventually decays over time but DMI leads to broader power curves. Overall, among the three configurations, the supercube consistently outperforms both open and closed chains in the XXZ model, mirroring its advantage in the Ising case.
Across both Hamiltonian models, the supercube configuration consistently delivers superior performance, combining high ergotropy, strong charging power, and stable charging cycles. In the Ising model, DMI enhances charging power for open and supercube chains but degrades performance in the closed chain. In contrast, under the XXZ model, DMI significantly improves the closed and supercube chains, with the latter displaying clean sinusoidal cycles and negligible residual energy, making it particularly attractive for rechargeable QB applications. Thus, while the influence of DMI depends on both model and topology, the supercube architecture remains the most robust and efficient design. Motivated by the superior performance of the supercube configuration across both Ising and XXZ models, we now examine how variations in interaction parameters and structural modifications affect its charging behaviour.
Varying interaction strengths in Supercube:
We study the effect of varying the DMI strength $D$ and the Heisenberg coupling constant $J$ on the charging behaviour of the supercube spin chain in XXZ model illustrated in figure LABEL:fig:erg_p_cube. Figure LABEL:fig-a:erg_vary_DM shows ergotropy for different values of $D$ keeping $J=1$ fixed. When $D\neq 1.7$, the optimal sinusoidal charging-discharging cycle disappears and the maximum ergotropy decreases, indicating that there exists an optimal DMI strength for sustaining high-efficiency periodic energy transfer between the battery and the charging field. Figure LABEL:fig-b:erg_vary_J shows ergotropy for different values of $J$ keeping $D=1.7$ fixed. The periodic charging cycle disappears for $J\neq 1$, and ergotropy of the system significantly diminishes, becoming nearly zero for $J>2$. These observations suggest that the combination $D=1.7$ and $J=1$ are optimal points for the supercube QB model under XXZ Hamiltonian. This behaviour further indicates the presence of a resonance-like condition, where the balance between exchange interaction ($J$) and spin–orbit coupling effects induced by DMI enables efficient cyclic energy transfer.
Adding diagonals in supercube configuration:
We investigate the effect of adding additional connections between the qubits in supercube configuration. In figure LABEL:fig-a:body_diag, we introduce interactions between the qubits located at the end points of two cross-body diagonals of the supercube. In figure LABEL:fig-b:face_diag we add interactions along the face diagonals on the top face of the cube. In both cases, we observed a noticeable decay in the amplitude of the charging-discharging cycle. When all four cross-body diagonals are added, the oscillation amplitude decreases further. Adding all face diagonals results in complete suppression of the sinusoidal nature of charging curve. Similarly, the sinusoidal nature vanishes when both body and face diagonals are introduced simultaneously into the system. These observations suggest that the addition of connections between qubits beyond the original supercube structure disrupts the energy exchange that supports optimal battery performance, and plays a detrimental role in charging of the supercube configuration. Hence, the original supercube model appears to represent a finely tuned structure to facilitate efficient charging cycles.
IV.2.4 Extending the battery to 12 qubits:
We extend the supercube architecture from 8 qubits to 12 qubits by adding an additional face to the original cube to investigate the effect of increasing the system size in this configuration. From the ergotropy plot shown in figure LABEL:fig-b:erg_12q, we observe that the system initially charges to a high-energy state which is slightly less than the ideal maximum of $2n\hbar\omega_{0}=24$. The subsequent discharging cycles show a decay in oscillation amplitude. Importantly, the 12-qubit system also relaxes to a state very close to the ground (passive) state after each cycle, leaving only a negligible amount of residual energy that cannot be extracted. This indicates that increasing the number of qubits does not disrupt the characteristic sinusoidal charging–discharging behavior. However, the cycle efficiency in the 12-qubit case is somewhat reduced compared to the 8-qubit supercube spin chain.
Beyond this supercube extension, we also explore alternative 12-qubit connectivities based on the cuboctahedron and icosahedron skeletons. The cuboctahedral structure yields an almost ideal sinusoidal curve. The cuboctahedral structure yields an almost ideal sinusoidal curve for $J=1$ and $D=1.94$, with minimal non-extractable residual energy, as illustrated in Fig. LABEL:fig-b:octa_erg. In the case of the icosahedron, we observe an even closer approximation to perfect charging–discharging cycles for $J=1$ and $D=2.06$, as shown in Fig. LABEL:fig-b:icosa_erg. Notably, this configuration also reaches the maximum ergotropy value more rapidly than the simple cubic extension.
V Conclusion
In this work, we studied the charging performance of eight-qubit Heisenberg spin chain quantum batteries (QBs) under the influence of Dzyaloshinskii-Moriya interaction (DMI). We compared three distinct spin chain configurations, namely, open boundary, closed boundary, and supercube, under both Ising and XXZ Hamiltonians. Our results show that the supercube consistently exhibits superior performance among all three configurations. For the Ising model, DMI enhances ergotropy in both the open boundary and supercube spin chains. In the XXZ model, the supercube configuration achieves a stable, near-ideal sinusoidal charging–discharging cycle, enabling periodic charging with negligible residual energy per cycle for optimal interaction strengths $D=1.7$ and $J=1$. We also investigated the effect of including the battery Hamiltonian $H_{B}$ in the charging process through the influence parameter $\lambda$. For all configurations and models, increasing $\lambda$ reduces the peak ergotropy. Interestingly, in the XXZ model, ergotropy for $D=1.7$ and $\lambda>0$ is higher than the $\lambda=0$ case for most of the evolution.
Additionally, we explored the effect of structural modifications of the supercube configuration. Introducing extra interactions between qubits disrupts the oscillatory behavior observed in the XXZ model. Extending the system size to 12 qubits preserves the sinusoidal nature of the ergotropy curve, but the system no longer discharges fully to its ground state, resulting in residual energy stored in the system. Notably, the 12-qubit icosahedral structure revives a perfect charging–discharging cycle, fully returning to the ground state without residual energy, thereby recovering the efficiency lost in the simple cubic extension. Although entanglement often influences QB performance, it has been shown that entanglement generation is not necessary for optimal work extraction [60]. In line with this, although our study does not explicitly examine entanglement, the results demonstrate enhancements in ergotropy and power without any direct connection to entanglement. Our findings provide insights into designing scalable and efficient quantum batteries and guide future exploration of optimized interaction topologies and system sizes. Interestingly, supercube-configured spin chain quantum batteries achieve high-efficiency, near-ideal charging cycles without requiring entanglement, providing a scalable route for next-generation quantum energy storage.
Code and data availability
The used codes and data sets are available in the GitHub repository. The code provided utilizes the QuantumToolbox package in both Python and Julia. [61, 62].